OpenStudy (yourgirlbri):

A 45 gram sample of a substance that's used to sterilize surgical instruments has a k-value of 0.15 Formula: N=N0e^-kt Find the substance's half-life, in days. Round to nearest tenth.

OpenStudy (photon336):

we can solve for t then plug in the numbers

OpenStudy (photon336):

\[N = N_{0}e^{-kt}\] \[\frac{ N }{ N_{0} } = e^{-kt}\] \[Ln(\frac{ N }{ N_{0} }) = Ln(e^{-kt})\] \[Ln(\frac{ N }{ N_{0} }) = -kt \] so this is how you would isolate t \[\frac{ Ln(\frac{ N }{ N_{0} }) }{ -k } = t \] now, all you need to do is to plug in your values. N0 would be 45 so N would be half of that 22.5

OpenStudy (photon336):

take a look at the math, you notice that you first solve for the variable you're looking for then plug in the numbers afterwards.

OpenStudy (yourgirlbri):

I'm not good at this stuff :/ I'm really confused

OpenStudy (photon336):

So here is the formula re-arranged for the variable that we want to find \[Ln(\frac{ N }{ N_{0} })*k^{-1} = t \]

OpenStudy (photon336):

\[N = our~final~ample \] \[N_{0} = our~initial~mass\]

OpenStudy (yourgirlbri):

I never actually learned this stuff properly so that's why I'm really confused, I just switched to online classes

OpenStudy (photon336):

do you know what half life is?

OpenStudy (yourgirlbri):

half of a number? (guess)

OpenStudy (photon336):

it is the time, it takes for half of our sample to go away

OpenStudy (yourgirlbri):

okay still confused

OpenStudy (yourgirlbri):

I have 20 more questions and it's 1 am fml

OpenStudy (photon336):

so let's say if I have 50 grams of a sample right? and it takes 20 minutes to get to 25 what can you conclude form this?

OpenStudy (yourgirlbri):

it decreaes?

OpenStudy (photon336):

yes but it decreases by half

OpenStudy (photon336):

so the time it takes for the sample to decrease by half is called the half life.

OpenStudy (yourgirlbri):

okay